# City-on-Web: Real-time Neural Rendering of Large-scale Scenes on the Web

### v1

@@ 7 Proof of 3D consistency
By integrating the rendering results of each block’s shader through the volume rendering between blocks, the final diffuse color $\boldsymbol{c}_{d}$ and specular feature $\boldsymbol{F}$ of the ray can be obtained.
$\boldsymbol{h}(\boldsymbol{r})=\sum_{k}\prod_{j=1}^{k-1}(1-\alpha^{j})\cdot\boldsymbol{h}^{k}$ (9)
Note that
$\quad\ 1-\alpha^{k}$ (10)
$=1-\sum_{i=1}^{N_{k}}\prod_{j=1}^{i-1}(1-\alpha_{j}^{k})\cdot\alpha_{i}^{k}$
@@
Assuming the block positions are already depth-sorted, meaning there are $M$ sampling points along the ray, with each block $k$ containing $N_{k}$ sampling points, let $\alpha_{i}$ denote the $i$-th sampling point on the ray, and $\alpha_{i}^{k}$ denote the $i$-th sampling point in block $k$ along the ray, then
$\boldsymbol{h}(\boldsymbol{r})$ $=\sum_{k}\prod_{j=1}^{k-1}\prod^{N_{j}}_{i=1}(1-\alpha_{i}^{j})\cdot\boldsymbol{h}^{k}$ (11)
@@
$=\sum_{k}\prod^{N_{1}+\cdots+N_{k-1}}_{i=1}(1-\alpha_{i}^{j})\cdot\left(\sum_{i=1}^{N_{k}}\prod_{j=1}^{i-1}(1-\alpha_{j}^{k})\cdot\alpha_{i}^{k}\boldsymbol{h}_{i}^{k}\right)$
@@
$=\sum_{k}\sum_{i=1}^{N_{k}}\prod^{N_{1}+\cdots+N_{k-1}}_{i=1}(1-\alpha_{i}^{j})\prod_{j=1}^{i-1}(1-\alpha_{j}^{k})\cdot\alpha_{i}^{k}\boldsymbol{h}_{i}^{k}$
$=\sum_{i=1}^{M}\prod^{i-1}_{i=1}(1-\alpha_{j})\alpha_{i}\boldsymbol{h}_{i}$
Eq. 11 shows that the diffuse color and specular feature we finally obtain by depth-sorting blocks and alpha blending along blocks are consistent with the results of volume rendering integration along the entire ray. Therefore, our method ensures the three-dimensional consistency of rendering.
@@
8 Dataset
To demonstrate the effectiveness of our method, experiments are conducted on a variety of large scenes. The main experiments reported in this paper involve three types of environmental scene datasets of various scales. The Campus dataset is captured at a altitude of about 180 meters, covering an area of approximately 960,000 $m^{2}$. The Matrix City Dataset, captured at a altitude of 200 meters, is sparser than the Campus dataset, thus covering a larger area. The Mill 19 dataset covers a total of about 200,000 $m^{2}$. We adopt a circular data capture method for photographing, as shown in Fig. 7. We find that this method often results in a higher overlap rate, allowing for a more accurate estimation of camera poses. Our dataset was captured over 8 hours on a cloudy day, with a fixed exposure setting to ensure almost identical appearance of photos taken at different times. We used Colmap to estimate camera poses. Feature matching was done using a vocabulary tree, followed by a hierarchical mapper followed by a few iterations of triangulation and bundle adjustment to estimate camera poses.
Figure 8: Visulization of ou efficient non-overlapping partition strategy. The red square represents the $[-1,1]^{2}$.
9 Efficiency of Non-overlapping Partition Strategy
In the stage of segmenting the scene into distinct blocks, we initially rotate the scene to align it parallel with the xy-plane, then proceed to segment the entire space based on the xy coordinates of spatial points. This strategy’s merit lies in ensuring each segmented block is a bounded area. This is in contrast to methods like segmentation strategy of block-nerf  [34] and  [37], which cannot assure boundedness in the reconstructed area, potentially leading to substantial memory resource wastage as illustrated in the  Fig. 8. Our approach allows us to represent the same size area with a bounded region of $[-1,1]^{3}$, maintaining the same representation resolution, instead of using an unbounded region that contract to $[-2,2]^{3}$ like in the MERF  [30] and Mip-NeRF 360 [1]. Consequently, this enables the reduction of the resolution of the xy-plane from $4096^{2}$ to $2048^{2}$ without loss in performance. Thus we can effectively reduce the usage of VRAM, especially across the three planes.
10 Impletation Details.
For blocks at the boundaries of the entire scene, an unbounded scene representation is required to represent areas outside the block boundaries. We follow the same approach as MERF to compute ray-AABB intersections trivially. To be specific, we employ the scene contraction function to project the scene external to the unit sphere into a cube, which has a radius of 2. The definition of the $j-th$ coordinate for a contracted point is as follows:
11 More Results
We provide additional qualitative results and on Matrix City, Mill 19, Campus dataset as shown in Fig. 9.
12 Discussion
Recently, some research has also enabled real-time rendering of large scenes. UE4-NERF [13], building on the MobileNerf framework, divides large scenes into smaller segments for reconstruction and then renders the large-scale scene using the mesh rasterization pipeline. Like MobileNerf, UE4-NERF begins with a $128^{3}$ grid, which assumes an even distribution of scene details in all directions. However, data captured through oblique photography often appears ’flat’, meaning there’s dense information when projected onto the xy plane, but sparser in the vertical direction due to mostly empty areas. Therefore, UE4-NERF needs more segments for large-scale reconstruction to ensure an even distribution of details in all directions in a block. For example, their Construction Site scene required about 40 blocks to reconstruct a $420m^{2}\times 240m^{2}$ area, leading to significant memory and VRAM usage, approximately 25GB. Even with UE4’s dynamic mesh-based loading, VRAM usage is around 11GB. Such high payload and VRAM demands make it challenging to extend this technology to web platforms and consumer-grade GPUs.
Additionally, NeuRas [3] has also made progress in real-time rendering of large scenes. It uses texture-less geometry from already reconstructed large scenes. NeuRas applys feature map textures and combines mesh rasterization results with view direction to query a small MLP for view-dependent colors. They optimize the feature texture and view-dependent MLP to enhance rendering results. However, this method also suffers from high memory usage. Our experiments show that obtaining texture-less geometry with ContextCapture [^1] https://www.bentley.com/software/itwin-capture-modeler/ for the Campus scene requires about 6GB, and surface reconstruction via neural rendering methods needs approximately 10GB without mesh simplification.
Both these methods have achieved excellent results in real-time rendering of large scenes, but their significant memory and VRAM requirements limit their expansion to web platforms. Our method can implement real-time rendering of large scenes with a payload and VRAM consumption acceptable for web platforms and consumer-grade graphics cards.
Figure 9: Qualitative visualization of geometry and rendering result on five datasets.

### v2

@@ A Derivation of Block-based Volume Rendering
By blending the rendering results of each block using Eq. (8), the final diffuse color $\boldsymbol{C}_{d}$ and specular feature $\boldsymbol{F}$ of the ray $\boldsymbol{r}$ can be obtained.
$\boldsymbol{C}_{d}(\boldsymbol{r})=\sum_{k=1}^{K}\prod_{j=1}^{k-1}(1-\alpha^{j})\cdot\boldsymbol{C}^{k}_{d}$ (12)
$\boldsymbol{F}(\boldsymbol{r})=\sum_{k=1}^{K}\prod_{j=1}^{k-1}(1-\alpha^{j})\cdot\boldsymbol{F}^{k}$
From Eq. 11, we get the following equation:
$\quad\ 1-\alpha^{k}$ $=1-\sum_{i=1}^{N_{k}}\prod_{j=1}^{i-1}(1-\alpha_{j}^{k})\cdot\alpha_{i}^{k}$ (13)
@@
Consequently, we can obtain the following by substituting Eq. 13 into Eq. 12.
$\boldsymbol{C}_{d}(\boldsymbol{r})$ $=\sum_{k=1}^{K}\prod_{j=1}^{k-1}(1-\alpha^{j})\cdot\boldsymbol{C}^{k}_{d}$ (14)
$=\sum_{k=1}^{K}\prod_{j=1}^{k-1}\prod^{N_{j}}_{i=1}(1-\alpha_{i}^{j})\cdot\boldsymbol{C}_{d}^{k}$
$=\sum_{k=1}^{K}\prod^{N_{1}+\cdots+N_{k-1}}_{i=1}(1-\alpha_{i})\cdot\boldsymbol{C}_{d}^{k}$
By substituting $\boldsymbol{C}_{d}^{k}$ from Eq. 11 into Eq. 14, we demonstrate the equivalence between Eq. 9 and Eq. 12.
$\boldsymbol{C}_{d}(\boldsymbol{r})$ $=\sum_{k=1}^{K}\prod^{N_{1}+\cdots+N_{k-1}}_{l=1}(1-\alpha_{l})\sum_{i=1}^{N_{k}}\prod_{j=1}^{i-1}(1-\alpha_{j}^{k})\cdot\alpha_{i}^{k}\boldsymbol{c}_{i}^{k}$
$=\underbrace{\alpha_{1}^{1}\boldsymbol{c}_{1}^{1}+\cdots+\prod_{i=1}^{N_{1}-1}(1-\alpha_{i}^{1})\alpha_{N_{1}}^{1}\boldsymbol{c}_{N_{1}}^{1}}_{block1}$
@@
$+\underbrace{\prod_{j=1}^{N_{1}}(1-\alpha_{j}^{1})\left(\alpha_{1}^{2}\boldsymbol{c}_{1}^{2}+\cdots+\prod_{i=1}^{N_{2}-1}(1-\alpha_{i}^{2})\alpha_{N_{2}}^{2}\boldsymbol{c}_{N_{2}}^{2}\right)}_{block2}+\cdots$
@@
$+\underbrace{\prod_{j=1}^{N_{1}}(1-\alpha_{j}^{1})\cdots\prod_{j=1}^{N_{K-1}}(1-\alpha_{j}^{K-1})\left(\alpha_{1}^{K}\boldsymbol{c}_{1}^{K}+\cdots+\prod_{i=1}^{N_{K}-1}(1-\alpha_{i}^{K})\alpha_{N_{K}}^{K}\boldsymbol{c}_{N_{K}}^{K}\right)}_{blockK}$
$=\underbrace{\alpha_{1}^{1}\boldsymbol{c}_{1}^{1}+\cdots+\prod_{i=1}^{N_{1}-1}(1-\alpha_{i}^{1})\alpha_{N_{1}}^{1}\boldsymbol{c}_{N_{1}}^{1}}_{block1}$ (15)
$+\underbrace{\prod_{j=1}^{N_{1}}(1-\alpha_{j}^{1})\alpha_{1}^{2}\boldsymbol{c}_{1}^{2}+\cdots+\prod_{j=1}^{N_{1}}(1-\alpha_{j}^{1})\prod_{i=1}^{N_{2}-1}(1-\alpha_{i}^{2})\alpha_{N_{2}}^{2}\boldsymbol{c}_{N_{2}}^{2}}_{block2}+\cdots$
$+\underbrace{\prod_{i=1}^{N_{1}}(1-\alpha_{i}^{1})\cdots\prod_{i=1}^{N_{K-1}}(1-\alpha_{i}^{K-1})\alpha_{1}^{K}\boldsymbol{c}_{1}^{K}+\cdots}_{blockK}\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad$
$+\underbrace{\prod_{i=1}^{N_{1}}(1-\alpha_{i}^{1})\cdots\prod_{i=1}^{N_{K-1}}(1-\alpha_{i}^{K-1})\prod_{i=1}^{N_{K}-1}(1-\alpha_{i}^{K})\alpha_{N_{K}}^{K}\boldsymbol{c}_{N_{K}}^{K})}_{blockK}$
$=\underbrace{\alpha_{1}\boldsymbol{c}_{1}+\cdots+\prod_{i=1}^{N_{1}-1}(1-\alpha_{i})\alpha_{N_{1}}\boldsymbol{c}_{N_{1}}}_{block1}$
$+\underbrace{\prod_{j=1}^{N_{1}}(1-\alpha_{j})\alpha_{N_{1}+1}\boldsymbol{c}_{N_{1}+1}+\cdots+\prod_{i=1}^{N_{1}+N_{2}-1}(1-\alpha_{i})\alpha_{N_{1}+N_{2}}\boldsymbol{c}_{N_{1}+N_{2}}}_{block2}+\cdots$
$+\underbrace{\prod_{j=1}^{N_{1}+\cdots+N_{K-1}}(1-\alpha_{j})\alpha_{N_{1}+\cdots+N_{K-1}+1}\boldsymbol{c}_{N_{1}+\cdots+N_{K-1}+1}+\cdots}_{blockK}$
$+\underbrace{\prod_{i=1}^{N_{1}+\cdots+N_{K}-1}(1-\alpha_{i})\alpha_{N_{1}+\cdots+N_{K}}\boldsymbol{c}_{N_{1}+\cdots+N_{K}}}_{blockK}$
$=\sum_{i=1}^{N_{1}+\cdots+N_{K}}\prod^{i-1}_{j=1}(1-\alpha_{j})\alpha_{i}\boldsymbol{c}_{i}$
$=\sum_{i=1}^{M}\prod^{i-1}_{j=1}(1-\alpha_{j})\alpha_{i}\boldsymbol{c}_{i}=\hat{\boldsymbol{C}}_{d}(\boldsymbol{r})$
It can be observed that block-based volume rendering approach in the resource-independent environment, such as in our case where multiple shaders are used to render the entire scene, is equivalent to the MERF’s volume rendering method [30] without deferred rendering.
B More Details on Experiments
In this section, we provide more details on our custom Campus dataset, implemention details and settings of comparison methods.
B.1 Dataset
The Campus dataset is captured at an altitude of about 180 meters, covering an area of approximately 960,000 $m^{2}$. We adopt a circular data capture method for areial photography, as shown in Fig. 6. We find that this method often results in a higher overlap rate, allowing for a more accurate estimation of camera poses. Our dataset was captured over 8 hours on a cloudy day, with a fixed exposure setting to ensure almost identical appearance of photos taken at different time. We used Colmap [31] to estimate camera poses. Feature matching was done using a vocabulary tree, followed by a hierarchical mapper followed by a few iterations of triangulation and bundle adjustment to estimate camera poses.
B.2 Implementation Details
For blocks at the boundaries of the entire scene, an unbounded scene representation is required to represent areas outside the block boundaries. We follow the same approach as MERF [30] to compute ray-AABB intersections trivially. To be specific, we employ the scene contraction function to project the scene external to the unit sphere into a cube, which has a radius of 2. The definition of the $j-th$ coordinate for a contracted point is as follows:
We use an A100 GPU for training. In Sections 5.1 and 5.2, we perform training for 50,000 iterations with a batch size of 16384 pixels for the Campus dataset, taking approximately 48 hours, while the training for other datasets takes around 24 hours. The experiments described in Sections 5.2 and 5.4 are subjected to training for 30,000 iterations with the same batch size, which is completed in approximately 12 hours. Training losses are initially balanced with $\lambda_{1}=1.0$, $\lambda_{2}=1.0$, $\lambda_{3}=0.01$, $\lambda_{4}=0.05$, $\lambda_{5}=0.001$ and sample $2^{14}$ samples for computing sparsity loss. During LOD genreation phase, we utilize the same loss function and hyperparameters as those employed during the training stage. We freeze the training of submodels for 5,000 iterations and then refine them and shared global deferred mlp jointly for an additional 10,000 iterations.
@@ B.2 Implementation Details
B.3 Comparative Method Settings
In Sec. 5.1, we adopt the official implementations of Mega-NeRF [39], NeRFacto [36], and Grid-NeRF [45]. Additionally, we use an unofficial implementation of Instant-NGP [^1] https://github.com/ashawkey/torch-ngp. Specifically, NeRFacto is trained with a batch size of 65,536 for 30,000 iterations. Instant-NGP [26] is trained for 500,000 iterations with a batch size of 4096. Grid-NeRF [45] is trained for 50,000 iterations with a batch size of 16384. Mega-NeRF [39] is trained for 500,000 iterations for each block, using a batch size of 1024. In Sec. 5.2, For MobileNeRF [5], We initialize a $192^{3}$ grid to generate polygonal meshes while adhering to default parameters for other setups. We use the open-source version [^2] https://github.com/hugoycj/torch-bakedsdf for BakedSDF [46], conducting training in two phases: 20,000 and 50,000 iterations, respectively with a batch size of 16,384 and use $1024^{3}$ to extract meshes by marching cubes [21]. We employ the official implementation of MERF [30] and 3D Gaussian Splatting [16]. During the MERF [30] experiments, the data is divided into four parts for a fair comparison with our method, and each block is trained with 32,768 batch size for 30,000 iterations, using default parameter settings. For 3D Gaussian splatting [16], we demonstrate the results after training for 30,000 and 100,000 iterations using the default settings.
C More Results
Figure 7: Qualitative comparison of real-time rendering methods.
We present more qualitative comparisons among our method, MERF [30] (4 blocks), and 3D Gaussian Splatting [16] (100k iters) on the Campus Dataset in Fig. 7. Additionally, we showcase rendering results from different LODs to demonstrate the appearance consistency across various LODs in Fig. 8.
Figure 8: Qualitative comparison of different levels-of-detail.
D Real-Time Viewer
Our real-time viewer platform is based on the MERF volume renderer [30], leveraging an OpenGL fragment shader to execute ray marching and deferred rendering. It accesses feature and density information through texture look-ups. However, we have implemented several improvements tailored for large scenes to enhance performance.
Dynamic Loading. We employ dynamic loading strategy to determine the level-of-detail and blocks to be loaded, thereby reducing VRAM usage. We use the center points obtained in Section 4.1 for each block on the $xy$ plane as the $xy$-components, and assign the height of the highest part within the block as the $z$-component. This defines the 3D center point for each block during the rendering stage. Additionally, we use the rectangle formed by the block’s four $xy$ corner points and the previously determined $z$ value to define the block’s region. By projecting this region onto the camera plane, we can determine whether the block is visible to the camera.
During the rendering phase, we first eliminate blocks that are not visible for cameras. Starting with the coarsest LOD, we check if the distance from the camera to all visible blocks within the finer LOD exceeds a certain threshold. For blocks beyond this threshold, rendering is performed using the resources loaded at the current LOD. Otherwise, the evaluation progresses to the subsequent, finer LOD. This stepwise process continues until we reach the finest LOD or complete the rendering of the entire image.
Depth sorting. We compute the distance from each block’s center point to the camera using only their $xy$-components, leveraging this calculated distance for depth sorting. We adopt this approach as our scene segmentation occurs solely in the $xy$ plane, without division along the $z$-axis.
E Discussion
Figure 9: Comparison to Mega-NeRF in Rubble dataset. The Rubble dataset exhibits significant lighting variations. The two images displayed above capture the same location but under vastly different lighting conditions. While our method reconstruct more details than Mega-NeRF [39], the deferred shading model has a limited capacity to represent the lighting and exposure variations as strong view-dependent effects. This limitations leads to marginally lower PSNR values.
Discussion. UE4-NeRF [10] and NeuRas [19], both mesh-based rendering approaches, offer fast rendering speed but face challenges when representing large, detail-rich scenes like those including dense foliage. These mesh representations can consume extensive memory, often amounting to several gigabytes. Additionally, 3D Gaussian Splatting [16], typically requires millions to tens of millions of points. The inherent properties of Gaussian splatting, which involve numerous Gauss-related attributes, further increase the VRAM needed for rendering. The substantial VRAM consumption characteristic of these methods poses significant challenges for implementing rendering on web platforms and resource-constrained devices, as they struggle to accommodate the high memory demands.
Limitation & Future Work. Our method still has some limitations. Since we derive alpha blending across shaders based on the Lambertian surface assumption, visible seams may occur at the boundaries between blocks on non-Lambertian surfaces, such as water surfaces. Combining physically-based rendering with multiple shaders blending may alleviate this problem. Additionally, while our approach achieves real-time rendering of large scenes on consumer-grade laptops, the inherently resource-intensive nature of large scenes makes that real-time rendering on mobile devices remains a challenge. Moreover, while our method recovers more intricate geometrical detail, it frequently results in color discrepancies with the ground truth image due to unstable lighting conditions and variable exposure, as shown in Fig. 9.